# Comparing Variable Selection and Model Averaging Methods for Logistic Regression

### v1

@@ Abstract
Model uncertainty is a central challenge in statistical models for binary outcomes such as logistic regression, arising when it is unclear which predictors should be included in the model. Many methods have been proposed to address this issue for logistic regression, but their relative performance under realistic conditions remains poorly understood. We therefore conducted a preregistered, simulation-based comparison of 28 established methods for variable selection and inference under model uncertainty, using 11 empirical datasets spanning a range of sample sizes and number of predictors, in cases both with and without separation. We found that Bayesian model averaging (BMA) methods based on gg–priors, particularly the $g=\max(n,p^{2})$, show the strongest overall performance when separation is absent. When separation occurs, penalized likelihood approaches, especially the LASSO, provide the most stable results, while BMA with the local empirical Bayes (EB-local) prior is competitive in both situations. These findings offer practical guidance for applied researchers on how to effectively address model uncertainty in logistic regression in modern empirical and machine learning research.
@@ Introduction
Model uncertainty is a central challenge in statistical inference, especially in regression settings where it is unclear in advance which predictors should be included in the model. In this context, each possible subset of the predictors can constitute a model, and even a modest number of predictors can give rise to a large number of possible models. Many methods have been proposed for addressing this problem, including both Bayesian and frequentist approaches. However, even though regression models are among the most widely used tools in applied statistics, the relative performance of these methods in realistic empirical settings had not been systematically evaluated. Porwal and Raftery [PorwalRaftery_2022] addressed this by conducting an extensive comparison of 21 commonly used methods for the linear regression model. They found that performance varied widely, and that three adaptive Bayesian model averaging approaches, which adapt to the analysis design and account for model uncertainty, performed best overall. Their results provide practical guidance for applied researchers and a benchmark for further methodological development.
Similar challenges arise when the outcome is binary and logistic regression is used. Logistic regression [Berkson1944, Cox1958, HosmerLemeshow2013] is a standard tool for modeling binary outcomes and is widely applied in fields such as epidemiology, social science, and machine learning. Here, we consider the logistic regression model
@@
As in the linear case, a major challenge is model uncertainty, which arises when there is a set of candidate predictors and it is unclear which ones to include in the model. In addition to uncertainty about variable inclusion, logistic regression introduces specific estimation difficulties. One such difficulty is separation, which occurs when a linear combination of predictors perfectly classifies the outcome. In these cases, unique maximum likelihood estimates may not exist [albert1984existence], leading to numerical instability and invalid inference. Separation is particularly common in small samples and high-dimensional settings, where complete or quasi-complete separation can arise even when the true signal is weak.
@@
Early applications of logistic regression typically relied on subject-matter expertise to determine which covariates to include in the model. This approach reflected the scientific priorities of the time, which emphasized interpretability and theoretical grounding, and was shaped by the computational constraints of early statistical analysis. Arguments and guidelines for doing this are described by [HosmerLemeshow2013, Bursac2008, Stoltzfus2011].
As logistic regression became more widely adopted and software implementations became more accessible, some analysts turned to a more inclusive strategy, fitting models with all available covariates. While this removed the need for subjective selection, it introduced new risks, including overfitting and degraded predictive performance. Stepwise selection procedures emerged as a practical alternative and gained popularity throughout the 1980s and 1990s [HosmerLemeshow2013]. These algorithms add or remove covariates based on statistical criteria such as likelihood ratio tests or information criteria. Even though stepwise methods have well-documented shortcomings [Derksen1992, miller2002subset, Freedman1983, Austin2004], especially when there is separation [heinze2006comparative, kotani2025goodness], they remain widely used to this day due to their simplicity and availability in standard software packages.
@@
Bayesian model averaging (BMA) is a widely studied framework for addressing model uncertainty in statistical inference. Instead of selecting a single model, BMA combines inferences across all plausible models, weighting each according to its posterior probability given the observed data. This approach offers a principled way to account for model uncertainty and has been shown to improve estimation and prediction in a variety of settings. Following its early development [Leamer1978, Raftery1988, George1993, Madigan1994], it has since been extensively applied to generalized linear models, including logistic regression [Tuchler01032008, Li02102018, ChenEtAl_1999, Wagner_2012, Raftery_1996_approx, Porwal_2024]. Several comprehensive reviews summarize the theoretical foundations and practical applications of BMA in regression [kass1995bayes, Hoeting1999, fernandez2001benchmark, Wasserman2000, Clyde2004, Fragoso2018, Forte2018, Kaplan_2021]. All of the Bayesian methods included in our comparison use BMA.
BMA has some good theoretical properties in general [raftery2003discussion]. BMA point estimators and predictions minimize mean squared error (MSE), and BMA estimation and prediction intervals are calibrated. BMA predictive distributions have optimal performance in the log-score sense [Madigan1994]. These properties hold on average over the prior distribution, extending similar results for Bayesian estimation [RubinSchenker1986], but the results are somewhat robust to this assumption [Mattei2020]. When the prior distribution is used as the distribution of parameter values over which performance is averaged, it has been referred to as the world distribution [Jeffreys1961], the practical distribution [raftery2003discussion], or the effect-size distribution [park2010estimation]. Analysis using this concept has been called “empirical frequentist” [Berger2021].
@@
Bayesian model averaging for logistic regression typically relies on adaptations of prior distributions originally developed for linear regression models. In particular, variants of Zellner’s $g$-prior [zellner1986assessing] and its extensions, including mixtures of $g$-priors, empirical Bayes adaptations [clyde2000flexible, george2000calibration], and fully Bayesian formulations such as the hyper-$g$ prior [liang2008mixtures], have been adapted for use in generalized linear models [Tuchler01032008, Li02102018]. These priors are widely implemented and support both variable selection and model averaging, although their theoretical justification is strongest under the Gaussian likelihood. In practice, they remain the default choice for BMA in logistic regression, and their comparative performance is best assessed empirically.
@@
All of the $g$-prior formulations evaluated in the linear setting are also available for logistic regression. This includes fixed-$g$ priors such as the unit information prior (with $g=n$), the benchmark prior (with $g=\max\{n,p^{2}\}$), and $g=\sqrt{n}$, which has been shown to perform well in high-dimensional settings [fernandez2001benchmark, young2014fast]. Other variants include the global and local empirical Bayes priors [HansenYu200] and the hyper-$g$ prior, which adapt $g$ based on the data through either empirical estimation or full Bayesian modeling. These priors differ in how they balance model complexity and parsimony, and they imply different effective prior sample sizes. The unit information prior has a close connection to the Bayesian Information Criterion (BIC), which can be used to approximate marginal likelihoods and posterior model probabilities in logistic regression [raftery1995bayesian]. Similarly, the Akaike Information Criterion (AIC) can be viewed as an approximation corresponding to $g=1$ [akaike1983information, burnham2002model], although recent work has shown that AIC-based selection may perform poorly under separation [kotani2025goodness]. Although $g$-priors are mathematically straightforward to extend to logistic models, the resulting posterior distributions are more complex due to the non-Gaussian likelihood. As a result, the implications of different choices of $g$ are less well understood in this setting. In the context of linear regression, Porwal and Raftery [PorwalRaftery_2022] found that the priors with $g=\sqrt{n}$, local empirical Bayes, and the hyper-$g$ prior performed best across a diverse set of real-data scenarios. Whether these results carry over to logistic regression is an important question that we investigate in this paper.
@@
In logistic regression, prior specification plays an especially important role in settings where separation is present. Proper prior distributions can mitigate this problem by regularizing extreme coefficient estimates and ensuring finite, stable inference. The use of proper but weakly informative priors is generally recommended to address separation and to improve robustness, interpretability, and computational stability in logistic regression [gelman2008prior, greenland2016penalization, discacciati2021weakly]. While priors such as the Cauchy and log-F distributions are not commonly used within BMA frameworks, their role as heavy-tailed regularizers is conceptually similar to that of robust and mixture $g$-priors. These priors offer varying degrees of shrinkage and regularization, and in practice, their ability to stabilize inference in the presence of separation may differ substantially.
Several alternative priors have been proposed and implemented in BMA software for logistic regression. For example, the BAS package offers the robust prior, which is a mixture of $g$-priors designed for outlier resistance; the intrinsic prior, which truncates the support of $g$ to downweight extreme values; and the Confluent Compound Hypergeometric (CCH) prior, which provides a flexible hierarchical structure on $g$ [forte2018methods]. These alternatives provide additional options for stabilizing estimation and model selection in difficult settings, although they have not been studied as extensively in the logistic case [Li02102018].
@@
In the frequentist setting, penalized likelihood approaches for logistic regression convert variable selection into an optimization problem. The objective typically involves the negative log-likelihood function with a penalty term $h_{\lambda}(\beta_{1},\dots,\beta_{)}$ on the coefficients $\bm{\beta}$. The penalized logistic regression estimate is
@@
The LASSO [Least Absolute Shrinkage and Selection Operator; tibshirani1996regression], where the penalty is $h_{\lambda}(\beta)=\lambda\sum_{j=1}^{p}|\beta_{j}|$, remains one of the most widely used methods. It encourages sparse solutions by shrinking some coefficients to exactly zero. Its popularity stems from its ability to perform variable selection while remaining computationally efficient in high-dimensional settings [buhlmann2011statistics]. For logistic regression, efficient algorithms based on coordinate descent have been developed to optimize the penalized log-likelihood, thus scaling up LASSO to larger problems [park2007l1, method_glmnet].
@@
However, the known limitations of LASSO in linear regression also apply to logistic regression. The oracle property ensures that the true predictors will be included in the selected model asymptotically, but not that all selected predictors are truly relevant, even asymptotically. As a result, LASSO can produce many false positives, even in large samples. It also tends to over-shrink large signals, which can lead to biased estimates [kyung2010penalized]. LASSO can also behave erratically in the presence of highly correlated predictors. It tends to select one variable from a group of collinear covariates and often discards the others, even if they carry important but weaker signals. This behavior can result in model instability and biased coefficient estimates, as relevant variables may be excluded due to their correlation with stronger predictors [holmes2011discussion].
@@
Several extensions have been proposed to address these concerns. The ridge penalty [hoerl1970ridge, lecessie1992ridge], $h_{\lambda}(\beta)=\lambda\sum_{j=1}^{p}\beta_{j}^{2}$, does not induce sparsity, but it improves stability and prediction when predictors are highly correlated. The elastic net combines LASSO and ridge penalties to encourage group selection and stability across correlated features [zou2005regularization]. Non-convex penalties such as SCAD [fan2001variable] and MCP [zhang2010nearly, breheny2011coordinate] reduce bias in large coefficients by applying weaker penalization at higher magnitudes. However, they are more difficult to optimize and are less commonly used in practice due to computational and tuning complexities. An alternative form of penalization is used in Firth’s bias-reduced logistic regression, which applies a penalty derived from Jeffreys’ invariant prior to reduce the first-order bias of the maximum likelihood estimator [firth1993bias]. This method yields finite estimates even in the presence of complete separation and has been stated to be particularly useful in small-sample settings [heinze2002solution].
@@
A common issue with penalized likelihood approaches is the lack of uncertainty quantification. In these approaches, variable selection is the outcome of a constrained optimization problem, and not a probabilistic statement of inclusion [Park2008, fouskakis2015power, womack2014inference]. Consequently, the zeros induced may not correspond to the variables that would be excluded under a full probabilistic model selection framework [hans2009bayesian]. These methods also provide no natural way to account for model uncertainty, which can lead to overconfident or unstable inference.
@@
Several methods have been proposed to address the lack of uncertainty quantification in penalized regression. One approach is the desparsified or debiased LASSO, which corrects the bias introduced by penalization and allows for asymptotically valid confidence intervals and hypothesis testing in high-dimensional settings [van2014asymptotically, javanmard2014confidence]. A related alternative is the Induced Smoothed LASSO (IS-LASSO) [method_islasso], which replaces the non-differentiable $\ell_{1}$ penalty with a smooth approximation. This enables likelihood-based inference, such as Wald and score tests, while maintaining sparsity. Bayesian approaches, such as EM Variable Selection (EMVS) [Rockova2014] and the spike-and-slab LASSO [rockova2018spike], combine ideas from model averaging and penalization to allow uncertainty quantification in principle. However, to our knowledge these ideas have not yet been implemented for logistic regression, and the ideas for uncertainty quantification are not yet fully supported in publicly available software.
@@
While this wide range of methods has been proposed for variable selection and inference in logistic regression, it remains unclear which ones perform best across the diverse settings encountered in practice. Penalized likelihood approaches such as the LASSO are widely used due to their computational efficiency and accessibility through software like glmnet [method_glmnet], while Bayesian methods offer a principled framework for uncertainty quantification but are less commonly adopted. Beyond the challenge of selecting relevant predictors, empirical analyses often encounter issues such as separation, particularly in small samples or high-dimensional settings. These practical complications can compromise inference and highlight the need for systematic comparisons across methods.
@@ Results
Fig. 1 and Fig. 2 summarize the results. Fig. 1 presents the results for datasets without separation, whereas Fig. 2 shows the results for datasets with separation. The tables report standardized performance metrics for point estimation (RMSE), interval estimation (MIS), prediction (Brier), and overall model selection (AUPRC). All metrics are standardized such that lower values indicate better performance and are ranked relative to the Spike–and–Slab method. Ranking scores were then calculated from these standardized metrics. The Full Score is defined as the average across all performance metrics, the Available Score as the average across the metrics available for each method, and the Partial Score as the average based on the RMSE and Brier score metrics, which are available for all methods. The methods are ranked according to the Partial Score. CPU time (in minutes) and failure rates are also reported, but were not used to compute the scores or determine the ranking. Complete simulation details are provided in the Materials and Methods, and per-metric and per-dataset results are available in the Supporting Information. In the Materials and Methods section, we detail the discrepancies between the preregistration plan and the final analyses. The most notable deviation is that we did not preregister monitoring for separation, nor the decision to present results stratified by the presence or absence of separation. This was not anticipated at the preregistration stage, as we did not foresee the extent to which separation would influence the analysis.
Across datasets without separation (Fig. 1), Bayesian model averaging (BMA) methods from the BAS package generally perform best. Among these, the so-called benchmark prior with $g=\max(n,p^{2})$ achieves the best scores overall, closely followed by the BIC, CCH, hyper–$g/n$, Beta-prime, and $g=\sqrt{n}$ priors. The Hyper-$g$, robust, intrinsic, and Spike and Slab priors also show strong and consistent performance across all metrics. The EB-local prior also performs competitively, though it slightly underperforms in prediction relative to the aforementioned priors. In contrast, EB-global performs poorly in both estimation and prediction, whereas the AIC and $g=4$ methods show weaker performance in estimation; all three are ranked below the reference method. All Bayesian methods exhibit a failure rate below 1%. The BMA method implemented in the BMA package shows slightly poorer predictive accuracy and exhibits relatively slow computational performance on average.
@@
Among the Bayesian approaches, the EB-local and Spike–and–Slab methods stand out due to their robust estimation and interval calibration. The EB-local method ranked ninth and the Spike-and-Slab method ranked tenth. However, the Spike-and-Slab method tends to be relatively slow. Several $g$-prior methods, as well as the AIC method, demonstrate significant performance degradation, particularly in point and interval estimation. However, some methods, such as $g=\max(n,p^{2})$, BIC, and BMA, maintain relatively strong performance in prediction and overall model selection, despite underperforming in estimation. All Bayesian methods had failure rates below 1%.

### v2

@@ Abstract
Model uncertainty is a central challenge in statistical models for binary outcomes such as logistic regression, arising when it is unclear which predictors should be included in the model. Many methods have been proposed to address this issue for logistic regression, but their relative performance under realistic conditions remains poorly understood. We therefore conducted a preregistered, simulation-based comparison of 28 established methods for variable selection and inference under model uncertainty, using 11 empirical datasets spanning a range of sample sizes and number of predictors, in cases both with and without separation. We found that Bayesian model averaging (BMA) methods based on gg–priors, particularly $g=\max(n,p^{2})$, show the strongest overall performance when separation is absent. When separation occurs, penalized likelihood approaches, especially the LASSO, provide the most stable results, while BMA with the local empirical Bayes (EB-local) prior is competitive in both situations. These findings offer practical guidance for applied researchers on how to effectively address model uncertainty in logistic regression in modern empirical and machine learning research.
@@ Introduction
Model uncertainty is a central challenge in statistical inference, especially in regression settings where it is unclear in advance which predictors should be included in the model. In this context, each possible subset of the predictors can constitute a model, and even a modest number of predictors can give rise to a large number of possible models. Many methods have been proposed for addressing this problem, including both Bayesian and frequentist approaches. However, even though regression models are among the most widely used tools in applied statistics, the relative performance of these methods in realistic empirical settings had not been systematically evaluated. Porwal and Raftery [81] addressed this by conducting an extensive comparison of 21 commonly used methods for the linear regression model. They found that performance varied widely, and that three adaptive Bayesian model averaging approaches, which adapt to the analysis design and account for model uncertainty, performed best overall. Their results provide practical guidance for applied researchers and a benchmark for further methodological development.
Similar challenges arise when the outcome is binary and logistic regression is used. Logistic regression [8, 24, 56] is a standard tool for modeling binary outcomes and is widely applied in fields such as epidemiology, social science, and machine learning. Here, we consider the logistic regression model
@@
As in the linear case, a major challenge is model uncertainty, which arises when there is a set of candidate predictors and it is unclear which ones to include in the model. In addition to uncertainty about variable inclusion, logistic regression introduces specific estimation difficulties. One such difficulty is separation, which occurs when a linear combination of predictors perfectly classifies the outcome. In these cases, unique maximum likelihood estimates may not exist [3], leading to numerical instability and invalid inference. Separation is particularly common in small samples and high-dimensional settings, where complete or quasi-complete separation can arise even when the true signal is weak.
@@
Early applications of logistic regression typically relied on subject-matter expertise to determine which covariates to include in the model. This approach reflected the scientific priorities of the time, which emphasized interpretability and theoretical grounding, and was shaped by the computational constraints of early statistical analysis. Arguments and guidelines for doing this are described by [56, 15, 97].
As logistic regression became more widely adopted and software implementations became more accessible, some analysts turned to a more inclusive strategy, fitting models with all available covariates. While this removed the need for subjective selection, it introduced new risks, including overfitting and degraded predictive performance. Stepwise selection procedures emerged as a practical alternative and gained popularity throughout the 1980s and 1990s [56]. These algorithms add or remove covariates based on statistical criteria such as likelihood ratio tests or information criteria. Even though stepwise methods have well-documented shortcomings [27, 77, 36, 4], especially when there is separation [50, 64], they remain widely used to this day due to their simplicity and availability in standard software packages.
@@
Bayesian model averaging (BMA) is a widely studied framework for addressing model uncertainty in statistical inference. Instead of selecting a single model, BMA combines inferences across all plausible models, weighting each according to its posterior probability given the observed data. This approach offers a principled way to account for model uncertainty and has been shown to improve estimation and prediction in a variety of settings. Following its early development [68, 86, 41, 74], it has since been extensively applied to generalized linear models, including logistic regression [101, 71, 19, 104, 85, 83]. Several comprehensive reviews summarize the theoretical foundations and practical applications of BMA in regression [62, 54, 30, 106, 23, 35, 32, 61]. All of the Bayesian methods included in our comparison use BMA.
BMA has some good theoretical properties in general [84]. BMA point estimators and predictions minimize mean squared error (MSE), and BMA estimation and prediction intervals are calibrated. BMA predictive distributions have optimal performance in the log-score sense [74]. These properties hold on average over the prior distribution, extending similar results for Bayesian estimation [90], but the results are somewhat robust to this assumption [76]. When the prior distribution is used as the distribution of parameter values over which performance is averaged, it has been referred to as the world distribution [60], the practical distribution [84], or the effect-size distribution [78]. Analysis using this concept has been called “empirical frequentist” [7].
@@
Bayesian model averaging for logistic regression typically relies on adaptations of prior distributions originally developed for linear regression models. In particular, variants of Zellner’s $g$-prior [111] and its extensions, including mixtures of $g$-priors, empirical Bayes adaptations [22, 40], and fully Bayesian formulations such as the hyper-$g$ prior [72], have been adapted for use in generalized linear models [101, 71]. These priors are widely implemented and support both variable selection and model averaging, although their theoretical justification is strongest under the Gaussian likelihood. In practice, they remain the default choice for BMA in logistic regression, and their comparative performance is best assessed empirically.
@@
All of the $g$-prior formulations evaluated in the linear setting are also available for logistic regression. This includes fixed-$g$ priors such as the unit information prior (with $g=n$), the $g=4$ prior [108], the benchmark prior (with $g=\max\{n,p^{2}\}$), and $g=\sqrt{n}$, which has been shown to perform well in high-dimensional settings [30, 110]. Other variants include the global and local empirical Bayes priors [40] and the hyper-$g$ prior, which adapt $g$ based on the data through either empirical estimation or full Bayesian modeling. These priors differ in how they balance model complexity and parsimony, and they imply different effective prior sample sizes. The unit information prior has a close connection to the Bayesian Information Criterion (BIC), which can be used to approximate marginal likelihoods and posterior model probabilities in logistic regression [87]. Similarly, the Akaike Information Criterion (AIC) can be viewed as an approximation corresponding to $g=1$ [2, 14], although recent work has shown that AIC-based selection may perform poorly under separation [64].
Several ways to extend $g$-priors to logistic regression models have been proposed [91, 71]. Here we focus on the proposal of Li and Clyde [71] to use a normal prior distribution with variance matrix proportional to the observed information matrix. Although $g$-priors can be extended to logistic regression models, the resulting posterior distributions are more complex due to the non-Gaussian likelihood. As a result, the implications of different choices of $g$ are less well understood in this setting. In the context of linear regression, Porwal and Raftery [81] found that the priors with $g=\sqrt{n}$, local empirical Bayes, and the hyper-$g$ prior performed best across a diverse set of real-data scenarios. Whether these results carry over to logistic regression is an important question that we investigate in this article.
@@
In logistic regression, prior specification plays an important role in settings where separation is present. Proper prior distributions can mitigate separation by regularizing extreme coefficient estimates and improving numerical stability. The use of proper but weakly informative priors is commonly recommended to address separation and improve robustness and interpretability, and heavy-tailed priors such as the Cauchy and the log-$F$ prior have been advocated for single-model inference [39, 46]. The Cauchy prior is closely related to the Zellner–Siow and mixture $g$-prior constructions often used in Bayesian model averaging. However, recent work shows that Cauchy-like tail behavior may yield ill-behaved posterior moments or extreme estimates under separation, motivating interest in lighter-tailed alternatives [42, 43]. Priors with different tail behavior offer varying degrees of shrinkage, and their ability to stabilize inference under separation can differ substantially.
Several extensions of the $g$-prior framework have been implemented in BMA software for logistic regression, including the BAS package. The robust prior was introduced to reduce sensitivity to the choice of a fixed $g$ [70, 69, 98]. Both the robust and intrinsic priors truncate the support of $g$ in a manner that depends on model dimension and sample size. As a result, their effective support can differ substantially from that of local empirical Bayes, benchmark, or hyper-$g/n$ priors when the sample size is large relative to the number of predictors [69, 71]. The Confluent Compound Hypergeometric (CCH) prior provides a unifying framework that encompasses these priors and their truncated variants as special cases [71]. Their behavior in logistic regression settings has received little systematic study.
@@
In the frequentist setting, penalized likelihood approaches for logistic regression convert variable selection into an optimization problem. The objective typically involves the negative log-likelihood function with a penalty term $h_{\lambda}(\beta_{1},\dots,\beta_{p})$ on the coefficients $\bm{\beta}$. The penalized logistic regression estimate is
@@
The LASSO [Least Absolute Shrinkage and Selection Operator; 100], where the penalty is $h_{\lambda}(\beta)=\lambda\sum_{j=1}^{p}|\beta_{j}|$, remains one of the most widely used methods. It encourages sparse solutions by shrinking some coefficients to exactly zero. Its popularity stems from its ability to perform variable selection while remaining computationally efficient in high-dimensional settings [13]. For logistic regression, efficient algorithms based on coordinate descent have been developed to optimize the penalized log-likelihood, thus scaling up LASSO to larger problems [79, 37].
@@
However, the known limitations of LASSO in linear regression also apply to logistic regression. The oracle property ensures that the true predictors will be included in the selected model asymptotically, but not that all selected predictors are truly relevant, even asymptotically. As a result, LASSO can produce many false positives, even in large samples. It also tends to over-shrink large signals, which can lead to biased estimates [17]. LASSO can also behave erratically in the presence of highly correlated predictors. It tends to select one variable from a group of collinear covariates and often discards the others, even if they carry important but weaker signals. This behavior can result in model instability and biased coefficient estimates, as relevant variables may be excluded due to their correlation with stronger predictors [55].
@@
Several extensions have been proposed to address these concerns. The ridge penalty [53, 67], $h_{\lambda}(\beta)=\lambda\sum_{j=1}^{p}\beta_{j}^{2}$, does not induce sparsity, but it improves stability and prediction when predictors are highly correlated. The elastic net combines LASSO and ridge penalties to encourage group selection and stability across correlated features [114]. Non-convex penalties such as SCAD [28] and MCP [113, 11] reduce bias in large coefficients by applying weaker penalization at higher magnitudes. However, they are more difficult to optimize and are less commonly used in practice due to computational and tuning complexities. An alternative form of penalization is used in Firth’s bias-reduced logistic regression, which applies a penalty derived from Jeffreys’ invariant prior to reduce the first-order bias of the maximum likelihood estimator [31]. This method yields finite estimates even in the presence of complete separation and has been stated to be particularly useful in small-sample settings [49].
@@
A common issue with penalized likelihood approaches is the lack of uncertainty quantification. In these approaches, variable selection is the outcome of a constrained optimization problem, and not a probabilistic statement of inclusion [80, 33, 107]. Consequently, the zeros induced may not correspond to the variables that would be excluded under a full probabilistic model selection framework [48]. These methods also provide no natural way to account for model uncertainty, which can lead to overconfident or unstable inference.
@@
Several methods have been proposed to address the lack of uncertainty quantification in penalized regression. One approach is the desparsified or debiased LASSO, which corrects the bias introduced by penalization and allows for asymptotically valid confidence intervals and hypothesis testing in high-dimensional settings [103, 59]. A related alternative is the Induced Smoothed LASSO (IS-LASSO) [20], which replaces the non-differentiable $\ell_{1}$ penalty with a smooth approximation. This enables likelihood-based inference, such as Wald and score tests, while maintaining sparsity. Bayesian approaches, such as EM Variable Selection (EMVS) [88] and the spike-and-slab LASSO [89], combine ideas from model averaging and penalization to allow uncertainty quantification in principle. However, to our knowledge these ideas have not yet been implemented for logistic regression, and the ideas for uncertainty quantification are not yet fully supported in publicly available software.
@@
While this wide range of methods has been proposed for variable selection and inference in logistic regression, it remains unclear which ones perform best across the diverse settings encountered in practice. Penalized likelihood approaches such as the LASSO are widely used due to their computational efficiency and accessibility through software like glmnet [37], while Bayesian methods offer a principled framework for uncertainty quantification but are less commonly adopted. Beyond the challenge of selecting relevant predictors, empirical analyses often encounter issues such as separation, particularly in small samples or high-dimensional settings. These practical complications can compromise inference and highlight the need for systematic comparisons across methods.
@@ Results
Fig. 1 and Fig. 2 summarize the results. Fig. 1 presents the results for datasets without separation, whereas Fig. 2 shows the results for datasets with separation. The tables report standardized performance metrics for point estimation (RMSE), interval estimation (IS), prediction (Brier), and overall model selection (AUPRC). All metrics are standardized such that lower values indicate better performance and are ranked relative to the Spike–and–Slab method. Ranking scores are then calculated from these standardized metrics. The Full Score is defined as the average across all performance metrics, the Available Score as the average across the metrics available for each method, and the Partial Score as the average based on the RMSE and Brier score metrics, which are available for all methods. The methods are ranked according to the Partial Score. CPU time (in minutes) and failure rates are also reported, but are not used to compute the scores or determine the ranking. Complete simulation details are provided in the Materials and Methods, and per-metric and per-dataset results are available in the Supporting Information. In the Materials and Methods section, we detail the discrepancies between the preregistration plan and the final analyses. The most notable deviation is that we did not preregister monitoring for separation, nor the decision to present results stratified by the presence or absence of separation. This was not anticipated at the preregistration stage, as we did not foresee the extent to which separation would influence the analysis.
Across datasets without separation (Fig. 1), Bayesian model averaging (BMA) methods from the BAS package generally perform best. Among these, the so-called benchmark prior with $g=\max(n,p^{2})$ achieves the best scores overall, closely followed by the BIC.BAS, CCH, hyper–$g/n$, Beta-prime, and $g=\sqrt{n}$ priors. The Hyper-$g$, robust, intrinsic, and Spike and Slab priors also show strong and consistent performance across all metrics. The EB-local prior also performs competitively, though it slightly underperforms in prediction relative to the aforementioned priors. In contrast, EB-global performs poorly in both estimation and prediction, whereas the AIC and $g=4$ methods show weaker performance in estimation; all three are ranked below the reference method. An additional examination of the posterior expected model size revealed that priors such as Hyper-$g$, EB.local, and $g=4$ tend to favor larger models, particularly in low-dimensional datasets, which partly explains their relatively weaker performance in estimation. In contrast, $g=\max(n,p^{2})$ and BIC.BAS consistently yielded more parsimonious models across all datasets. All Bayesian methods exhibit a failure rate below 1%. The BIC.BMA method implemented in the BMA package shows slightly poorer predictive accuracy and exhibits relatively slow computational performance on average.
@@
Among the Bayesian approaches, the EB-local and Spike–and–Slab methods stand out due to their robust estimation and interval calibration. The EB-local method ranked ninth and the Spike-and-Slab method ranked tenth. However, the Spike-and-Slab method tends to be relatively slow. Several $g$-prior methods, as well as the AIC method, demonstrate significant performance degradation, particularly in point and interval estimation. However, some methods, such as $g=\max(n,p^{2})$, BIC.BAS, and BIC.BMA, maintain relatively strong performance in prediction and overall model selection, despite underperforming in estimation. All Bayesian methods had failure rates below 1%.